# {cutout}03.2cm55cm1 empty PDF-WuKong         : A Large Multimodal Model for Efficient Long PDF Reading with End-to-End Sparse Sampling

### v2

@@ Abstract
Multimodal document understanding is a challenging task to process and comprehend large amounts of textual and visual information. Recent advances in Large Language Models (LLMs) have significantly improved the performance of this task. However, existing methods typically focus on either plain text or a limited number of document images, struggling to handle long PDF documents with interleaved text and images, especially for academic papers. In this paper, we introduce PDF-WuKong, a multimodal large language model (MLLM) which is designed to enhance multimodal question-answering (QA) for long PDF documents. PDF-WuKong incorporates a sparse sampler that operates on both text and image representations, significantly improving the efficiency and capability of the MLLM. The sparse sampler is integrated with the MLLM’s image encoder and selects the paragraphs or diagrams most pertinent to user queries for processing by the language model. To effectively train and evaluate our model, we construct PaperPDF, a dataset consisting of a broad collection of English and Chinese academic papers. Multiple strategies are proposed to automatically generate 1.1 million QA pairs along with their corresponding evidence sources. Experimental results demonstrate the superiority and high efficiency of our approach over other models on the task of long multimodal document understanding, surpassing proprietary products by an average of 8.6% on F1. Our code and dataset will be released at https://github.com/yh-hust/PDF-Wukong.
footnotetext: Equal contribution; Corresponding author
@@ 1 Introduction
The advent of Large Language Models (LLMs) has significantly advanced the field of PDF document understanding [1, 2], where these models have demonstrated impressive capabilities in processing and generating human-like text. However, they still face many challenges when it comes to lengthy PDF documents with interlaced text and images, such as academic papers.
@@
To handle lengthy documents, current research in multimodal document understanding with LLMs primarily follows two mainstream technical routes. The first route is based on pure text modality understanding. As shown in Fig. 1(a), these approaches typically consider the parsed OCR results from PDF documents, and convert all visual elements into textual representations (e.g., captions and OCR content extracted from figures and tables). They then employ long-context LLMs [3, 4, 5] or utilize Retrieval Augmented Generation (RAG) techniques [6, 7, 8, 9] to process the textual content. The main disadvantage of this approach is the significant loss of visual information inherent in multimodal documents, making it challenging to support answering fine-grained visual-related questions.
@@
Figure 1: Method comparison for long multi-page PDF document understanding. (a) Plain text solution: long-context/RAG LLMs for parsed pure text content. (b) Pure vision solution: VDU models for page-level encoding and feature interaction. (c) Our method is based on end-to-end sparse sampling for long PDFs with interleaved text and images.
@@
The second route is based on pure visual modality understanding. As illustrated in Fig.1(b), this approach generally avoids parsing PDF documents and instead treats each page as an image, and utilizing multimodal LLMs for document understanding tasks[10, 11, 12]. However, high-resolution images and numerous pages generate a large number of visual tokens, significantly increasing the models’ input token length and leading to scalability issues. This makes it challenging to efficiently process multi-page long documents. While some recent multi-page VDU models handle up to 8 pages [13] or 20 pages [14], they typically encode each page separately and perform page-level visual feature interactions such as concatenation [15, 16]. Nevertheless, as the number of pages grows, computational resource consumption still escalates substantially, rendering these models inefficient for processing longer documents.
@@
Considering the limitations of existing methods in multimodal understanding of long PDF documents, we propose a new MLLM architecture with end-to-end sparse sampling, named PDF-WuKong. Since most user queries are only related to a small part of the content in a long document, the sparse sampling can significantly remove redundant noise information. It encodes text paragraphs and diagrams in the parsed PDF document, utilizing both text and image representations to identify and extract the most relevant evidence in response to a user’s query. The sampled sparse evidence significantly reduces the number of input tokens of LLM and this process is independent of the length of the input documents. Moreover, the sparse sampler and LLM can be integrated in an end-to-end manner for training and inference, optimizing the performance of multimodal representation and question answering while improving time efficiency. It is worth noting that this sparse sampler is a plug-and-play design that can be applied to any MLLMs. Another important characteristic is that it can naturally provide strong interpretability for the question answering.
@@
We train PDF-WuKong on PaperPDF dataset, complemented by general-domain document question-answering datasets. Experimental results substantiate the effectiveness and efficiency of our approach to the task of long multimodal PDF understanding. PDF-WuKong significantly outperforms potential open-source models that may be applied to this task. It also surpasses some proprietary products for document understanding on our proposed PaperPDF benchmark. As the number of document pages increases, its accuracy and efficiency will not decrease significantly. It also achieves competitive performance on several document-oriented VQA datasets, especially multi-page benchmarks like DUDE [17]. Besides, for the recent benchmark MM-NIAH [18] of long multimodal documents, PDF-WuKong also outperforms other models with fewer parameters. Our model achieves the best performance on multimodal content with a context length of 64K.
@@
  • We introduce a large multimodal model for long PDF understanding with end-to-end sparse sampling, achieving accurate and efficient PDF question answering.
@@
Input modality Type Number of tokens Models
Plain text Long-context Linear increase LongLoRA [4], LongLLaMA [5], YaRN [3]
RAG w/o Linear increase Graph RAG [9], DISC-LawLLM [6], RAPTOR [19]
Pure vision Single-page w/o Linear increase UniDoc [20], DocOwl [21], Vary [12], UReader [22], TextMonkey [10], LLaVA-NeXT [23], XC2-4KHD [24], InternVL-V1.5 [11]
Multi-page Linear increase Hi-VT5 [14], GRAM [16], Fox [13], DocOwl2 [25], CREAM [26]
Text and images Unlimited-page w/o Linear increase PDF-WuKong (Ours)
@@ 2.1 Document Understanding Datasets
Early datasets focused on NLP tasks like summarization [27] and QA [28] of plain text, while visual document datasets targeted text perception tasks such as Document Layout Analysis (DLA) [29, 30, 31] and Key Information Extraction (KIE) [32, 33, 34]. Recent multimodal document QA datasets include DocVQA [35] and OCRVQA [36] for single-page documents, ChartQA [37] and ChartX [38] focusing on visual reasoning in charts. Datasets like ArXivQA [39] and InfoVQA [40] enhance MLLMs’ abilities on academic and infographic documents. However, these datasets are limited to single-page tasks, and current MLLMs [41, 11] perform well on them.
Multi-page QA datasets like MP-DocVQA [14], DUDE [17], DocGenome [42], and MM-NIAH [18] require understanding content relationships via multi-hop reasoning. Yet, answers in these datasets lack evidence and reliable interpretability, especially for questions needing multiple pieces of evidence from long documents.
@@ 2.2 Document Understanding Methods
Existing methods focus on plain text or limited document images. Text-based approaches convert documents into plain text using OCR and then employ long-context mechanisms like sparse attention [4], memory networks [5], or position interpolation [3]. Retrieval-augmented generation methods [6, 19, 9] also handle long texts effectively. These approaches struggle with fine-grained visual understanding.
Another solution, visual document understanding, treats each page as an image. MLLMs like UniDoc [20], mPLUG-DocOwl [21], and Vary [12] perform OCR-free understanding. Models such as UReader [22] and TextMonkey [10] divide high-resolution pages into patches. InternLM-XC2-4KHD [24] and InternVL-V1.5 [11] introduce a dynamic resolution mechanism with automatic patch configuration. However, reliance on high resolution increases token counts and isn’t scalable to multi-page documents.
For multi-page documents, models like Hi-VT5 [14], GRAM [16], Fox [13], CREAM [26] and mPLUG-DocOwl2 [25] encode pages separately and perform page-level interactions. More pages generate more visual tokens, increasing resource consumption and inefficiency for longer documents. Thus, we propose parsing documents into interleaved text and images, followed by sparse sampling in an end-to-end manner. Tab. 1 summarizes these methods.
@@ 3 Methodology
Figure 2: The overall structure of PDF-WuKong consists of a document parser, a sparse sampler and a large language model.

### v3

# Equal contribution Project lead

@@ Abstract
Multimodal document understanding is a challenging task to process and comprehend large amounts of textual and visual information. Recent advances in Large Language Models (LLMs) have significantly improved the performance of this task. However, existing methods typically focus on either plain text or a limited number of document images, struggling to handle long PDF documents with interleaved text and images, especially for academic papers. In this paper, we introduce PDF-WuKong, a multimodal large language model (MLLM) that is designed to enhance multimodal question-answering (QA) for long PDF documents. PDF-WuKong incorporates a sparse sampler that operates on both text and image representations, significantly improving the efficiency and capability of the MLLM. The sparse sampler selects the paragraphs or diagrams most pertinent to user queries. To effectively train and evaluate our model, we construct PaperPDF, a dataset consisting of a broad collection of English and Chinese academic papers. Multiple strategies are proposed to build high-quality 1.1 million QA pairs along with their corresponding evidence sources. Experimental results demonstrate the superiority and high efficiency of our approach over other models on the task of long multimodal document understanding, surpassing proprietary products by an average of 8.6% on F1. Our code and dataset will be released at https://github.com/yh-hust/PDF-Wukong.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has significantly advanced the field of PDF document understanding saad2023pdftriage ; jacob2024constructing ; zhang2025smaller , where these models have demonstrated impressive capabilities in processing and generating human-like text. However, they still face many challenges when it comes to lengthy PDF documents with interlaced text and images, such as academic papers.
@@
To handle lengthy documents, current research in multimodal document understanding with LLMs primarily follows two mainstream technical routes. The first route is based on pure text modality understanding. As shown in Fig. 1(a), these approaches typically consider the parsed OCR results from PDF documents, and convert all visual elements into textual representations (e.g., captions and OCR content extracted from figures and tables). They then employ long-context LLMs peng2024yarn ; chen2024longlora ; NEURIPS2023_8511d06d or utilize Text Retrieval Augmented Generation (RAG) techniques yue2023disc ; tang2024multihop ; edge2024local to process the textual content. The main disadvantage of this approach is the significant loss of visual information inherent in multimodal documents, making it challenging to support answering fine-grained visual-related questions.
@@
Figure 1: Method comparison for long multi-page PDF document understanding. (a) Plain text solution: long-context LLMs and TextRAG-based LLMs for parsed pure text content. (b) Pure vision solution: Multi-page MLLMs with page-level feature interaction and VisualRAG-based MLLMs for page images. (c) Our method is based on end-to-end multimodal sparse sampling for long PDFs with interleaved text and images. Note: “LLM” in (c) is the language backbone component integrated within MLLM.
@@
The second route is based on pure visual modality understanding. As illustrated in Fig.1(b), this approach generally avoids parsing PDF documents and instead treats each page as an image, and utilizing multimodal LLMs for document understanding tasks liu2024textmonkey ; chen2024far ; wei2025vary . However, high-resolution images and numerous pages generate a large number of visual tokens, significantly increasing the models’ input token length and leading to scalability issues. This makes it challenging to process multi-page long documents efficiently. While some recent multi-page visual document understanding models handle up to 8 pages liu2024focus or 20 pages TITO2023109834 , they typically encode each page separately and perform page-level visual feature interactions such as concatenation SlideVQA2023 ; blau2024gram . Nevertheless, as the number of pages grows, computational resource consumption escalates substantially, rendering these models inefficient for processing longer documents. Although some recent VisualRAG-based models yu2024visrag ; chen2025svrag try to find document pages related to the query, they perform poorly for text-intensive scenarios and cannot provide more fine-grained evidence.
@@
Considering the limitations of existing methods in multimodal understanding of long PDF documents, we propose a new MLLM architecture with end-to-end sparse sampling, named PDF-WuKong. Since most user queries are only related to a small part of the content in a long document, the sparse sampling can significantly remove redundant noise information. It encodes text paragraphs and diagrams in the parsed PDF document, utilizing both text and image representations to identify and extract the most relevant evidence in response to a user’s query. The sampled sparse evidence significantly reduces the number of input tokens of the LLM and this process is independent of the length of the input documents. Moreover, the sparse sampler and LLM can be integrated in an end-to-end manner for training and inference, optimizing the performance of multimodal representation and question answering while improving time efficiency. It is worth noting that this sparse sampler is a flexible and generic design that can be applied to various MLLM architectures through fine-tuning. Another important characteristic is that it can naturally provide strong interpretability for question answering.
@@
Input modality Type Number of tokens Models
Plain text Long-context Linear increase LongLoRA chen2024longlora , LongLLaMA NEURIPS2023_8511d06d , YaRN peng2024yarn
Text RAG w/o Linear increase Graph RAG edge2024local , DISC-LawLLM yue2023disc , RAPTOR sarthi2024raptor
Pure vision Multi-page Linear increase Fox liu2024focus , DocOwl2 hu2024mplugdocowl2 , MiniCPM-V yao2024minicpm ; yu2025minicpm , InternVL InternVL2.5 ; wang2025internvl3_5 , Qwen-VL Qwen2.5-VL ; Qwen3-VL , Granite Vision team2025granite , DocSeeker yan2026docseeker
Visual RAG w/o Linear increase VisRAG yu2024visrag , SV-RAG chen2025svrag
Text and images Multi-page Linear increase Hi-VT5 TITO2023109834 , GRAM blau2024gram
Multimodal RAG w/o Linear increase PDF-WuKong (Ours)
We train PDF-WuKong on the PaperPDF dataset, complemented by general-domain document question-answering datasets. Experimental results substantiate the effectiveness and efficiency of our approach to the task of long multimodal PDF understanding. PDF-WuKong significantly outperforms potential open-source models that may be applied to this task. It also surpasses some proprietary products for document understanding on our proposed PaperPDF benchmark. As the number of document pages increases, its accuracy and efficiency will not decrease significantly. It also achieves competitive performance on several document-oriented VQA datasets, especially multi-page benchmarks like DUDE van2023document . Besides, for the recent benchmark MM-NIAH wang2024needle of long multimodal documents, PDF-WuKong also outperforms other models with fewer parameters. Our model achieves the best performance on multimodal content with a context length of 64K.
@@
  • We introduce a large multimodal model for long PDF understanding with end-to-end sparse sampling, achieving accurate and efficient performance and providing explainable evidence for the answer.
@@ 2.1 Document Understanding Datasets
Earlier document understanding datasets only focused on the NLP tasks such as summarization huang2021efficient , and QA dasigi2021dataset of plain text. Meanwhile, there were several visual document datasets mainly aimed at text perception tasks such as Document Layout Analysis (DLA) zhong2019publaynet ; li2020docbank ; pfitzmann2022doclaynet and Key Information Extraction (KIE) simsa2023docile ; park2019cord ; huang2019icdar2019 . Recently, more datasets have been proposed for multimodal document QA across various scenarios. For example, DocVQA Mathew_2021_WACV and OCRVQA  MishraSSC19 provide single-page QA data on books and business documents. ChartQA masry2022chartqa and ChartX xia2024chartx focus on visual reasoning for chart documents. ArXivQA li2024multimodal extracts scientific figure-caption data from the arxiv papers to enhance the academic ability of MLLMs. InfoVQA Mathew_2022_WACV contains many infographic documents that are a combination of textual, graphical and visual elements. However, these visual document datasets only define the single-page VQA task.
There are also some multi-page QA datasets SlideVQA2023 ; van2023document ; TITO2023109834 that require the model to understand the content relationship via multi-hop reasoning and capture the crucial information from multi-page documents. MP-DocVQA TITO2023109834 extends DocVQA Mathew_2021_WACV by adding the context pages. DUDE van2023document constructs multi-page QA data from multi-industry and multi-domain documents. Additionally, a large-scale dataset designed for the comprehensive understanding of lecture slides has been introduced zhang2025towards . DocGenome xia2024docgenome was constructed as a scientific document benchmark. MM-NIAH wang2024needle is a benchmark evaluating the capability of MLLMs to comprehend long multimodal documents, which requires the model to answer according to the key information scattered throughout the document. However, the answers in these datasets lack evidence and cannot provide reliable interpretability, especially for questions that require referring to multiple pieces of evidence in long documents.
@@ 2.2 Document Understanding Methods
Existing document understanding methods typically focus on either plain text or a limited number of document images. Methods that rely on pure text modality aim to process documents by first converting them into plain text through document parsing or Optical Character Recognition (OCR) zhao2024cbnet ; zheng2024cdistnet ; wu2024end ; liu2025multi . They then employ efficient long-context mechanisms to handle long texts, such as sparse attention chen2024longlora , memory networks NEURIPS2023_8511d06d , or position interpolation peng2024yarn . Besides, the methods based on retrieval-augmented generation (RAG) yue2023disc ; sarthi2024raptor ; edge2024local also show impressive capabilities for long texts. While these approaches can integrate visual elements by including image captions or transforming images into natural language descriptions, they struggle with a fine-grained understanding of visual information. This restricts their effectiveness in tasks requiring detailed interpretation of textual and visual components within documents.
Another solution is visual document understanding with purely visual input, treating each document page as an image. Many MLLMs such as UniDoc feng2023unidocuniversallargemultimodal , mPLUG-DocOwl ye2023mplugdocowl , StrucTexTv3 lyu2024structextv3 and Vary wei2025vary can perform this task in an OCR-free manner. Other researchers have advanced the understanding of high-resolution document pages by dividing input images into smaller patches, such as UReader ye2023ureader , TextMonkey liu2024textmonkey , and LLaVA-NeXT liu2024llavanext . DocPedia feng2023docpedia processes high-resolution images in the frequency domain via the DCT transformation. InternVL-V1.5 chen2024far introduces a dynamic resolution mechanism to capture more details. The reliance on high resolution results in a higher number of tokens and cannot be extended to multi-page documents. Recently, CoPali faysse2025colpali uses Vision LLM to retrieve query-related image patches in a single-page image, but it has not been applied to multi-page documents and only conducts the retrieval task without the QA task.
There are several visual models specifically designed for multi-page documents. Fox liu2024focus unifies all image tokens of up to 8 pages into a sequence to achieve multi-page QA. mPLUG-DocOwl2 hu2024mplugdocowl2 compresses each high-resolution document image into 324 tokens, guided by low-resolution global visual features. Notably, DocSeeker yan2026docseeker embeds the retrieval capability directly into the model rather than relying on a separate RAG pipeline, introducing structured visual reasoning to achieve accurate and fine-grained evidence grounding in multi-page documents. To improve model interpretability, DocThinker yu2025docthinker leverages rule-based reinforcement learning for explainable document understanding. Besides, some general MLLMs such as MiniCPM-V yao2024minicpm ; yu2025minicpm , InternVL InternVL2.5 ; wang2025internvl3_5 , Qwen-VL Qwen2.5-VL ; Qwen3-VL , and Granite Vision team2025granite can already support multi-image input by training with multi-image samples. These models encode each page image separately and then perform page-level visual feature interactions such as concatenation. However, more pages generate more visual tokens, which brings greater resource consumption to the LLM. This makes them inefficient and unable to process longer documents.
Recently, some document understanding MLLMs based on pure visual RAG yu2024visrag ; chen2025svrag ; yinmstar have been proposed. They usually separate the page image retrieval and document understanding generation into two independent stages during training and inference. They still treat each page as a separate image and use pure visual retrieval to obtain page-level evidence. These retrieval schemes remain challenging for text-intensive document understanding due to insufficient text recognition capabilities. Moreover, they can only obtain page-level evidence rather than more fine-grained evidence.
Beyond plain text and pure vision approaches, certain specialized models for multi-page documents attempt to integrate multimodal information by concatenating OCR tokens with visual features for each page. Hi-VT5 TITO2023109834 and GRAM blau2024gram are representative works in this category. They combine tokens of multiple pages through hierarchical transformer architecture TITO2023109834 or global-local reasoning blau2024gram , being able to handle up to 20 pages. However, their token count still increases linearly with the number of pages.
Therefore, we propose first to parse the lengthy document into interleaved text and image content and then perform multimodal sparse sampling in an end-to-end manner. Tab. 1 summarizes the attributes of each type of method.
@@ 3 Methodology
Figure 2: The overall structure of PDF-WuKong consists of a document parser, a multimodal sparse sampler that integrates end-to-end with an LLM. “LLM” in the figure refers to the language backbone within the MLLM.
3.1 Motivation and Method Overview
The central motivation behind PDF-WuKong is to empower multimodal large language models (MLLMs) to efficiently deal with sophisticated multimodal reasoning and comprehension tasks on lengthy multimodal PDF documents, particularly academic papers densely populated with paragraphs, figures, and tables. Directly inputting entire documents, including all textual paragraphs, figures, and tables, into existing MLLMs is computationally prohibitive due to their substantial length. Furthermore, crucial information relevant to a user’s query may be diluted by irrelevant content, making it challenging for the model to accurately locate and reason about key evidence.
To address these challenges, our overarching insight is to leverage sparse multimodal sampling to selectively identify and retrieve the most relevant text and image blocks given a user query. Crucially, instead of introducing a fully independent multimodal retrieval model, we propose a lightweight retrieval strategy that reuses the visual encoder from the pretrained MLLM. This strategy reduces computational redundancy and facilitates efficient embedding caching, enabling fast inference and effective end-to-end training.
As illustrated in Fig. 2, PDF-WuKong comprises three main components: 1) Document Parsing converts PDF documents into structured, interleaved sequences of text and image blocks (Sec. 3.2); 2) Multimodal Sparse Sampling encodes and caches compact embeddings for retrieval, efficiently identifying query-relevant sparse subsets of blocks (Sec. 3.3); 3) Multimodal Response Generation takes the retrieved sparse set of multimodal blocks and synthesizes concise, accurate responses to user queries using a pretrained MLLM (Sec. 3.4). The overall inference process is summarized in Algorithm 1.
3.2 Document Parsing and Representation
Given a PDF document $D$, we first parse it into a structured sequence of multimodal blocks, explicitly preserving its layout and reading order. Specifically, we extract:
- A set of text blocks $\{T_{1},T_{2},\dots,T_{n}\}$, each typically corresponding to a paragraph.
- A set of image blocks $\{I_{1},I_{2},\dots,I_{m}\}$, each including figures or tables with their captions.
If an image block has an accompanying textual caption, we explicitly associate them together. The structured multimodal content is stored as an XML file and parsed offline using open-source PDF parsing tools Grobid GROBID and MinerU wang2024mineruopensourcesolutionprecise .
To ensure robust parsing of complex academic layouts, we employ a hybrid parsing pipeline combining Grobid GROBID for metadata extraction and MinerU wang2024mineruopensourcesolutionprecise for fine-grained layout analysis. We selected this specific combination after comparing multiple open-source alternatives. First, both tools are lightweight and computationally efficient, enabling the scalable processing of our massive 1.1M document corpus. Second, MinerU is specifically optimized for academic documents; it demonstrates superior performance in parsing complex elements such as mathematical formulas and charts compared to traditional rule-based parsers. Crucially, its robust support for Chinese character recognition ensures high-quality parsing for the Chinese subset of our bilingual dataset. This pipeline effectively mitigates common parsing errors in mixed-content documents.
3.3 Multimodal Sparse Sampling
To effectively handle lengthy documents, our sparse sampling approach aims to identify a concise subset of the most query-relevant text and image blocks. To achieve this, we encode each block into a compact embedding vector, enabling efficient retrieval through similarity matching in a unified embedding space.
Text Embedding. We introduce a lightweight text encoder $E_{T}(\cdot)$ for each text block $T_{i}$, resulting in a single text embedding vector $e_{T_{i}}\in\mathbb{R}^{d_{T}}$ that concisely captures semantic content.
Image Embedding. For each image block $I_{j}$, we reuse the visual encoder $E_{V}(\cdot)$ from the pretrained MLLM to obtain the visual sequence embeddings $e_{I_{j}}\in\mathbb{R}^{n\times d_{V}}$, where $n$ is the number of visual tokens and $d_{v}$ is the visual embedding dimension. To unify visual and textual embeddings into a shared embedding space, we aggregate the visual sequence into a single compact vector and project it into the textual embedding space $d_{t}$ using the projection head $MLP_{proj}$}:
$e_{I_{j}}^{\prime}=MLP_{proj}(MeanPool(e_{I_{j}})),\quad e_{I_{j}}^{\prime}\in\mathbb{R}^{d_{T}}.$} (1)
If an image block has an associated textual caption, we further enrich its embedding by encoding the caption text using the same text encoder $E_{T}(\cdot)$ and obtain a single caption embedding vector. The final embedding $e_{I_{j}}^{\text{fused}}$ for the image is obtained by element-wise addition of the projected visual embedding and the caption embedding.
Query-driven Retrieval. Given a user query $q$, we also encode it using the text encoder $E_{T}(\cdot)$ to obtain a compact query embedding vector in $e_{q}\in\mathbb{R}^{d_{T}}$. Then, we can compute cosine similarity between the query embedding and all cached compact embeddings of text blocks ($\{e_{T_{i}}\}$) and image blocks ($\{e_{I_{j}}^{\text{fused}}\}$). The top-$K$ most similar blocks are selected as sparse evidence for multimodal response generation.
Offline Caching. All compact embeddings for text and image blocks, along with the original visual sequence embeddings, are computed offline and cached. This offline caching strategy significantly speeds up inference by avoiding redundant visual encoding computations during response generation.
3.4 Multimodal Response Generation
Given the retrieved sparse multimodal blocks and the user query, we generate the final response through the pretrained MLLM. For image blocks, we reuse the cached visual sequence embeddings $\{e_{I_{j}}\}$, containing all image patch tokens. Since these are derived from the MLLM’s native visual encoder, they can be directly projected into the input space of the MLLM’s language backbone using the original connection module $MLP_{conn}$, preserving fine-grained visual details:
$\hat{e_{I_{j}}}=MLP_{conn}(e_{I_{j}}),\quad\hat{e_{I_{j}}}\in\mathbb{R}^{n\times d_{llm}},$} (2)
$\hat{e_{I_{j}}}$ serves as the visual context tokens fed into the LLM backbone. This approach eliminates redundant visual encoding operations at inference time.
Algorithm 1 Inference pipeline for PDF-WuKong
1:Input: PDF document $D$, user query $q$
2:Output: Generated answer $a$
3:Initialize: Text encoder $E_{T}$, image encoder $E_{V}$, LLM, and two MLP projectors $\text{MLP}_{\text{proj}}$, $\text{MLP}_{\text{conn}}$
4:Stage 1: Document Parsing
5:Parse the input document $D$ into text blocks and images with their captions:
$\{T_{i}\},\{(I_{j},C_{j})\}\leftarrow\text{Parser}(D)$
6:Stage 2: Sparse Sampling
$\{e_{T_{i}}\}\leftarrow E_{T}(\{T_{i}\}),$
$\{e_{I_{j}}\}\leftarrow E_{V}(\{I_{j}\}),$
$\{e_{C_{j}}\}\leftarrow E_{T}(\{C_{j}\}),$
$\{e_{I_{j}}^{\prime}\}\leftarrow\text{MLP}_{\text{proj}}(MeanPool(\{e_{I_{j}}\})),$}
$\{e_{I_{j}}^{\text{fused}}\}\leftarrow\{e_{I_{j}}^{\prime}\}+\{e_{C_{j}}\}$
8:Encode the user query $q$:
$e_{q}\leftarrow E_{T}(q)$
9:Calculate the similarity between query embedding $e_{q}$ and cached text/image embeddings $\{e_{T_{i}}\}$, $\{e_{I_{j}}^{\text{fused}}\}$:
$S_{T}=\{\text{sim}(e_{q},e_{T_{i}})\mid i=1,2,\dots,n\},$
$S_{I}=\{\text{sim}(e_{q},e_{I_{j}}^{\text{fused}})\mid j=1,2,\dots,m\}$
10:Select top-$k$ relevant text blocks and images with captions:
$(T,I,C)_{top}\leftarrow\text{TopK}(S_{T},S_{I},k)$}
11:Stage 3: Answer Generation
12:Input the query $q$ and the selected top-$k$ text blocks $T_{i}$ and visual embeddings $e_{I_{j}}$ into the LLM:
$\{\hat{e_{I_{j}}}\}\leftarrow\text{MLP}_{\text{conn}}(\{e_{I_{j}}\}),$
$a\leftarrow\text{LLM}(q,(T_{i},\hat{e_{I_{j}}},C_{i})_{top})$}
13:Return the generated answer $a$.
Conversely, for text blocks, we provide raw textual input. Reusing retrieval embeddings is not feasible because 1) the text encoder (BGE-M3) and the MLLM’s language module use different tokenizers and vocabularies, and 2) LLMs are optimized to process discrete token indices mapped to their own internal embedding spaces. Providing raw text ensures the model utilizes its pre-trained linguistic capabilities to the fullest extent.
It is worth noting that we utilize the original visual embeddings $\{e_{I_{j}}\}$ for generation rather than the fused embeddings $\{e_{I_{j}}^{fused}\}$ used in retrieval. This decision is based on the fact that the $MLP_{conn}$ is pre-trained to align pure visual features with the LLM space; introducing fused embeddings at this stage would cause a distribution mismatch.
To encourage transparent reasoning and enhance semantic understanding, we explicitly annotate each input block with metadata such as block indices and section headings. Specifically, for retrieved image blocks, their corresponding captions are also provided as plain text inputs to the LLM alongside the visual embeddings. The MLLM is trained to explicitly reference relevant block indices when generating responses, thereby enhancing interpretability and evidence grounding.
3.5 Joint End-to-End Training Objectives
PDF-WuKong is trained end-to-end using a combination of two complementary learning objectives: (1) a contrastive representation learning loss $\mathcal{L}_{\text{CTR}}$ for effective embedding alignment, and (2) a cross-entropy loss $\mathcal{L}_{\text{CE}}$ for accurate multimodal response generation.
Contrastive Representation Loss ($\mathcal{L}_{\text{CTR}}$). For each training instance (query-answer pair), positive blocks ($B^{+}$) are those explicitly annotated as relevant evidence, while negative blocks ($B^{-}$) are randomly sampled irrelevant blocks from the document. Given a query embedding $e_{q}\in$, we optimize embeddings by minimizing the following contrastive loss:
$\mathcal{L}_{\text{CTR}}=-\frac{1}{|B^{+}|}\sum_{e^{+}\in B^{+}}\log\frac{\exp\bigl(\text{sim}(e_{q},e^{+})/\tau\bigr)}{\mathcal{Z}},$ (3)
where the normalization term $\mathcal{Z}$ sums over positive and negative samples:
$\mathcal{Z}=\exp\bigl(\text{sim}(e_{q},e^{+})/\tau\bigr)+\sum_{e^{-}\in B^{-}}\exp\bigl(\text{sim}(e_{q},e^{-})/\tau\bigr).$
Here, $\text{sim}(\cdot,\cdot)$ denotes cosine similarity, and $\tau$ is a temperature hyperparameter controlling the softness of the similarity distribution.
Cross-Entropy Response Generation Loss ($\mathcal{L}_{\text{CE}}$). The second objective trains the MLLM to generate accurate textual responses conditioned on the retrieved multimodal evidence and the query:
$\mathcal{L}_{\text{CE}}=-\sum_{l=1}^{L}\log p_{\theta}(r_{l}^{*}|q,B^{+},r_{<l}^{*}),$ (4)
where $r^{*}$ is the ground-truth response sequence, and $p_{\theta}$ is the MLLM.
Joint Optimization. Our end-to-end training objective integrates both losses:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CTR}}+\mathcal{L}_{\text{CE}}.$ (5)
Overall, PDF-WuKong introduces joint alignment to reconcile efficiency and fidelity: the visual features first map to a text-dominated metric space for coarse cross-modal retrieval, then to the LLM-aligned compositional space for fine-grained reasoning. This design mirrors human coarse-to-fine comprehension, filtering salient evidence via unified semantics before synthesizing grounded responses. Meanwhile, it avoids redundant recomputation through parameter-efficient adaptation and offline caching. The duality resolves modality granularity conflicts, ensuring computational tractability without sacrificing pretrained multimodal synergies.
4 PaperPDF Dataset